3x + 7 = (9 + x) + 2x
Can it be?
0 = -2

Answer by CharlesG2(834)

"3x + 7 = (9 + x) + 2x
Can it be?
0 = -2"

3x + 7 = (9 + x) + 2x
3x + 7 = 9 + x + 2x (removing parentheses)
3x + 7 = 9 + 3x (adding the x's on the right)
3x - 3x = 9 - 7 (bringing 3x's to left, and the constants to the right)
0 = 2, result on right is not -2 it is 2, and no 0 can not equal 2 or anything else other than 0
this problem has no solution
you may notice that 3x + 7 and 3x + 9 have the same coefficient on the x
which is 3
that means that in the form y = mx + b, where m is the slope, and b is the
y-intercept (vertical-intercept), that y = 3x + 7 and y = 3x + 9 are parallel lines, parallel lines never intersect, therefore no solution